OpenStudy (sleepyjess):

Last parabola question of the night :) Find the equation Focus at (-3, 4); directrix is the line y = 2

OpenStudy (sleepyjess):

I think once I find the vertex, I'll be fine, but I'm not sure how to find the vertex...

OpenStudy (sleepyjess):

@amistre64 , wanna help on one more? :P

OpenStudy (amistre64):

where does the vertex usually go?

OpenStudy (sleepyjess):

I know it's between the focus and directrix

OpenStudy (amistre64):

then 0 one of them out ...

OpenStudy (amistre64):

if y=2, then y-2 = 2-2 subtract 2 from all the y parts

OpenStudy (sleepyjess):

so, (3, 2)?

OpenStudy (sleepyjess):

*-3

OpenStudy (amistre64):

Focus at (-3, 4); directrix is the line y = 2 -2 -2 ----- ----- -3,0 0 so, we have a vertex that is between -3 and 0 (in this adjustment, just add back the 2 to move it back)

OpenStudy (amistre64):

ack ... i cant add worth nothing tonight -3,2 and y=0

OpenStudy (amistre64):